# Not All Attention is Needed: Parameter and Computation Efficient Tuning for Multi-modal Large Language Models via Effective Attention Skipping

### v2

@@ 4.3 Attention Redundancy Evaluation
In EAS, we adopt a $k$-arm bandit based algorithm to automatically evaluate the redundancy of MHAs in MLLMs for downstream task.
Concretely, given a MLLM, we first randomly sample and tune the subnetworks with PIAs as adapters for a few epochs. Afterwards, we initialize a numerical action vector, denoted as $\mathbf{s}\in\mathbb{R}^{n}$, and then we keep training and testing the MLLM with PIAs. In each evaluation step, we skip $k$ MHAs according to $\mathbf{s}$, based on which the action policy at the $t$ step is obtained by
@@
where $U(a,b)$ is the uniform distribution between $a$ and $b$. In this way, the effective modules are more likely to obtain a higher preference score. And then we skip $k$ MHA modules with the lowest preference scores:
@@
Then, we update the action preference of each MHA through comparisons. For each $T$ step, we sample $m$ distinct subnetworks, i.e, $\Phi^{(t)}=\{\rho^{(t)}_{1},\rho^{(t)}_{2},...,\rho^{(t)}_{m}\}$, according to Eq.14 and Eq.15. Based on the loss values $l^{(t)}_{i}$, the reward of a subnetwork is defined as $e^{-l^{(t)}_{i}}$, which is proportional to subnetwork performance.
According to the rewards, the action preference of MHA modules can be updated by
$s^{(t)}_{i}=$ $s^{(t-1)}_{i}+(\frac{1}{m}\sum_{h=1}^{m}e^{-l^{(t)}_{h}}-e^{-l^{(t)}_{j}})\pi^{(t)}_{i}(1-\pi^{(t)}_{i}),$ (16)
$\mathrm{where}\ \ i\in\rho^{(t)}_{j},\rho^{(t)}_{j}\in\Phi^{(t)}.$
@@
When the reward of the sampled subnetwork is larger than the average one $\frac{1}{m}\sum_{j=1}^{m}e^{-l^{(t)}_{j}}$, it suggests that the skipped MHA modules are less important. The change of their action preference in this step can be ontained by $\frac{1}{m}\sum_{j=1}^{m}e^{-l^{(t)}_{j}}-e^{-l^{(t)}_{i}}$. To smooth optimization, we introduce $\pi^{(t)}_{i}(1-\pi^{(t)}_{i})$ to control the learning rate.
Via a large number of trails, most subnetworks can be well evaluated. Finally, according to the action preference $\mathbf{s}$, which is used to reflect the redundancy, we can select that the most redundant MHA modules to skip, thereby reducing the computation complexity of MLLMs.
@@ 5.1 Datasets and Metrics
ScienceQA [36] is a multi-modal dataset for science question answering. This dataset is divided into three subsets, i.e., train, val, and test, which have $12,726$, $4,241$, and $4,241$ examples, respectively. The questions can be categorized into Natural Science (NAT), Social Science (SOC), Language Science (LAN), Text Context (TXT), Image Context (IMG), No Context (NO), Grades 1-6 (G1-6) and Grades 7-12 (G7-12), respectively. VQA2.0 [13] has $204,721$ images, each accompanied by at least three open-ended questions. We report the test-dev results from the online evaluation  [^4] https://eval.ai/web/challenges/challenge-page/830/overview. NLVR2 [45] is tailored for classifying triplets consisting of two images and a question, and it has $107,292$ examples of human-written sentences grounded with pairs of images. Flickr30K [43] is a cross-modal retrieval dataset, comprising $31,000$ images collected from Flickr.
@@ 5.2 Implementation Details
The base model employed in our study is LaVIN [38], a recently introduced MLLM. Following its default settings, we use LLaMA-7B [50] as the main network, and adopt ViT-L [9] as the image encoder. In terms of PIA, we set the hidden dimension to 32. Besides, we employ PIAs for parameter-efficient adaptations, similar to [38, 54]. These PIAs have a hidden dimension of 8, and only one path for adaptation is used, i.e., $f_{d1}$ in Eq. 4. We insert these PIAs before each attention module. For redundancy evaluation, we first train the subnetworks with randomly skipped modules for 5 epochs. Then we introduce redundancy evaluation in the next 5 epochs, where 3 subnetworks are sampled every 10 training steps. In addition, we also validate the proposed EAS on a conventional VL pre-trained model named METER [10]. Its attention modules in a co-attention layer are considered as independent potential skipped components. The PIAs for skipping attention modules are configured with a hidden dimension of 144, with the remaining settings following the defaults in DAS [53].
@@ 5.3.1 Quantitative Analysis
Comparison with DAS. We first compare our EAS with DAS [53] on LaVIN and METER on four VL benchmarks, of which results are given in Tab. 1 and Tab. 2. We first observe that both DAS and EAS can achieve competitive performance, while retaining parameter and computation efficiency. Speciffically, in Tab. 1, DAS can skip $4$ Transformer layers of LaVIN for better efficiency while retraining comparable performance, i.e $-0.26$ accuracy while $+14.8\%$ efficiency. Meanwhile, EAS-7B2 can improve the average accuracy and inference speed by $+0.62$ and 1.52×1.52$\times$1.52 × under the same parameter scale. With a little extra parameter cost, our EAS can skip $12$ MHAs while achieving better performance than both DAS and LaVIN. Notably, compared with LaVIN and DAS, EAS-7B12 can improve the inference speeds by $2.18$ and $1.90$ times, respectively. In Tab.2, we compare to DAS on three classic VL benchmarks, i.e. VQA, NLVR and Flickr30K. Compared with DAS, our EAS can update $23.8\%$ fewer parameters and improve inference by $+33.44\%$ with only $-0.54\%$ performance drop on average. Overall, these results well validate EAS towards the target of PCETL for MLLMs.
Table 1: Comparison among EAS and other LLM-Based methods on ScienceQA. The corner marks of PCETL methods represent the number of skipped modules. The best and second best results are marked in bold and underline, respectively.
Method Updated Parameters Inference Time [^3] Tested on ScienceQA with one A100 GPU. Subject Context Modality Grade Average
NAT SOC LAN TXT IMG NO G1-6 G7-12
Fully tuned methods
LLaVA-7B [32] 100.00% 2.19s 90.36 95.95 88.00 89.49 88.00 90.66 90.93 90.90 90.92
Parameter-efficient methods
LLaMA-Adapter [54] 0.03% 4.35s 84.37 88.30 84.36 83.72 80.32 86.90 85.83 84.05 85.19
LaVIN-7B [38] 0.06% 3.70s 89.25 94.94 85.24 88.51 87.46 88.08 90.16 88.07 89.41
LaVIN-7B-LoRA [17] 0.08% 2.22s 84.06 77.95 85.73 82.84 74.12 88.64 83.77 82.20 83.21
Parameter and computation efficient methods
DAS-7B4 [53] 0.07% 3.23s 88.99 94.60 85.09 87.88 86.51 88.36 89.72 88.13 89.15
EAS-7B2 (Ours) 0.07% 2.13s 89.48 93.93 87.00 88.32 87.36 89.76 90.20 88.99 89.77
EAS-7B12 (Ours) 0.11% 1.69s 89.52 95.61 86.36 88.71 88.25 88.85 91.26 87.67 89.98
Method Updated Parameters VQA NLVR2 Flickr30K Average
test-dev +FLOPs test-P +FLOPs IR/TR R@1 +FLOPs Acc. +FLOPs
Full Tuning 323.31M 77.43 +0.00 83.05 +0.00 82.22/94.30 +0.00 84.25 +0.00
Classifier Only - 69.93 +0.00 73.23 +0.00 78.80/89.00 +0.00 77.74 +0.00
Shallow Prompt [31] 0.30M 68.51 +28.71G 65.69 +26.84G 74.20/88.60 +28.71G 74.25 +28.71G
Deep Prompt [19] 1.84M 70.78 +6.53G 72.64 +5.59G 78.84/89.40 +6.53G 77.92 +6.53G
LoRA [17] 0.29M 74.00 +0.00 78.82 +0.00 79.86/92.60 +0.00 81.32 +0.00
Adapter [48] 5.34M 74.70 +1.64G 79.93 +1.38G 80.38/91.90 +1.64G 81.73 +1.64G
Scaled PA [15] 3.59M 75.11 +1.12G 80.38 +0.66G 80.40/93.20 +1.12G 82.27 +1.12G
DAS4 [53] 5.34M 74.80 -11.16G 80.11 -5.13G 80.12/91.80 -11.16G 81.71 -9.15G
EAS8(Ours) 4.07M 74.85 -10.65G 80.13 -7.38G 80.38/92.20 -9.12G 81.89 -9.05G
EAS10(Ours) 4.07M 74.82 -14.53G 80.08 -10.19G 79.78/90.40 -11.90G 81.27 -12.21G
@@
Comparison with the state-of-the-arts. In Tab. 1 and Tab. 2, we also compare EAS with more recent advancements of MLLMs and PETL methods. We can first observe that PETL methods achieve competitive or even better performance than the fully tuned methods, while retaining a very small number of parameters to update. However, the use of adapters also obviously slows down their inference speeds. For instance, the inference of LLaMA-Adapter in Tab. 1 consumes more than $4.35$ seconds for each example. In contrast, LoRA does not incur any additional overhead during inference, thus only taking $2.22$ seconds per example. Compared with these approaches, the results of EAS are more significant in both performance and efficiency. EAS-7B12 improve $+8.14\%$ performance of LoRA while improve inference speed up to 1.30×1.30$\times$1.30 × on ScienceQA. Meanwhile, in Tab. 2, we can observe that, EAS8 achieves $99.53\%$ performance of Scaled PA while decreasing $-12.28\%$ FLOPs. Overall, these results well validate the effectiveness of EAS towards the target of PCETL for MLLMs.
Table 3: Comparison of different redundancy evaluation candidates for EAS on ScienceQA.
@@
Candidate Modules Number of Skipped modules Updated Parameters Inference Time Average Performance
MHA 4 0.08% 2.08s 89.44
8 0.10% 1.91s 89.79
12 0.11% 1.69s 89.98
FFN 4 0.08% 2.26s 88.30
8 0.11% 2.18s 84.96
12 0.13% 2.11s 82.13
MHA or FFN 4 0.08% 2.13s 88.40
8 0.10% 2.09s 89.18
12 0.12% 1.78s 88.63
@@
Table 4: Ablation of Propagate-Information Adapter (PIA) on EAS-7B12. Here, Base refers to using only $f_{d1}$ in Eq. 4.
Setting Updated Parameters Inference Time Average Performace
Base 0.09% 2.94s 87.08
+ information exchange in Eq.4 0.11% 3.23s 90.05
+ frozen bias term in Eq.11 0.11% 2.97s 89.98
+ re-parameterzation 0.11% 1.69s 89.98
Table 5: Ablation of the skipping number.
Methods Updated Parameters Inference Time Average Performance
LaVIN-7B 0.06% 3.70s 89.41
EAS-7B4 0.08% 2.08s 89.44
EAS-7B8 0.10% 1.92s 89.79
EAS-7B12 0.11% 1.69s 89.98
EAS-7B16 0.13% 1.56s 88.42
Figure 4: The predictions of EAS-B12 and LaVIN-7B on ScienceQA. The accurate explanations for the answer are highlight in green, while the logically incorrect ones in red.
@@
Ablation Study. To examine the different designs of PIA, we further conduct extensive ablations in Tab.3-5. In Tab.3, we first ablate the choice of evaluation candidates for EAS on ScienceQA, including “MHA”, “FFN” and “MHA or FFN”. From this table, we can first see that skipping MHA is the best choice among three candidates. Notably, EAS can skip up to 12 MHAs with even better performance, while its efficiency is also much better than the others. In terms of FFN, its removal leads to slower inference speed than that of MHA. More importantly, when dropping more than 8 FFNs, the performance declines significantly, e.g., $-6.99\%$ by skipping 12 FFNs. This result well confirms our argument about the roles of MHA and FFN in MLLMs. In terms of “MHA or FFN”, this candidate is suboptimal for EAS, which can obtain a good trade-off between performance and efficiency, but it is still worse than “MHA”. Overall, these results well validate the motivation of EAS, i.e., not all attention is needed for MLLMs.
Tab. 4 ablates the designs of PIA. Here, “Base” refers to using only one path in the downsampling, i.e., only $f_{d1}$ in Eq. 4. We can see that without the two-path design for information exchange, its performance drops greatly, i.e., $-3.30\%$ than “+information exchange”. In order to realize re-parameterization, we freeze $avg(f_{d2}(\mathbf{X}^{(i)}))$ in Eq. 11, i.e., frozen bias term in Tab. 4. It can be seen that this setting barely affects performance, i.e., $-0.08\%$. In this case, EAS can conduct re-parameterization without declining performance, while improving inference speed greatly. Overall, Tab. 4 well confirms the design of information exchange and validates the re-parameterization in PIA.

### v3

@@ 4.3 Attention Redundancy Evaluation
In EAS, we adopt a $k$-arm bandit based algorithm to automatically evaluate the redundancy of MHAs in MLLMs for downstream task. This formulation addresses the combinatorial challenge of selecting $k$ redundant modules from $n$ MHAs. The search space of $\binom{n}{c}$ subnetworks is computationally hard to evaluate exhaustively. The k-armed bandit framework efficiently balances exploration of novel configurations with exploitation of high-performing subnetworks while providing regret-bounded optimization guarantees [auer2002finite]. In this way, we can obtain the accelerated model with the best performance at a given computational overhead.
Concretely, given a MLLM, we first randomly sample and tune the subnetworks with PIAs as adapters for a few epochs. Afterwards, we initialize a numerical action preference, denoted as $\mathbf{s}\in\mathbb{R}^{n}$, and then we keep training and testing the MLLM with PIAs. Every few steps, we will sample several substructures for comparison and update the action preference. In the comparison, we skip $k$ MHAs according to $\mathbf{s}$, and the action policy at the $t$ step is obtained by
@@
where $U(a,b)$ is the uniform distribution between $a$ and $b$. In this way, the effective modules are more likely to obtain a higher preference score. The action preference score $s_{i}^{(t)}$ directly quantifies the importance of the $i$-th MHA module. Lower scores indicate higher redundancy. For each subnetwork, we skip $k$ MHA modules with the lowest action policies:
@@
Then, we update the action preference of each MHA through comparisons. For each $T$ step, we first sample the action policies for these $m$ subnetworks, e.g., the action policy for the $j$-th subnetwork $\mathbf{\Pi}^{(t)}_{j}=[\pi^{(t)}_{j,1},\pi^{(t)}_{j,2},...,\pi^{(t)}_{j,n}$], according to Eq.14. Based on this, we can gain the skipped MHAs in these $m$ subnetworks, i.e, $\Phi^{(t)}=\{\rho^{(t)}_{1},\rho^{(t)}_{2},...,\rho^{(t)}_{m}\}$, according to Eq.15. And the skipped MHAs in the $j$-th subnetwork can be represent as $\rho^{(t)}_{j}=\{\mu_{j,1},\mu_{j,2},...,\mu_{j,m}\}$. Based on the loss values $l^{(t)}_{j}$ of the $j$-th subnetwork, the reward of it is defined as $e^{-l^{(t)}_{j}}$, which is proportional to subnetwork performance. According to the rewards, the action preference of MHA modules can be updated by
$s^{(t)}_{i}$= $s^{(t-1)}_{i}+(\frac{1}{m}\sum_{h=1}^{m}r^{(t)}_{h}-r^{(t)}_{j})\pi^{(t)}_{j,i}(1-\pi^{(t)}_{j,i})$, (16)
$\mathrm{where}\ \ i\notin\rho^{(t)}_{j},\rho^{(t)}_{j}\in\Phi^{(t)}.$
@@
When the reward of the sampled subnetwork is larger than the average one $\frac{1}{m}\sum_{j=1}^{m}e^{-l^{(t)}_{j}}$, it suggests that the activated MHA modules are more important. The change of their action preference in this step can be ontained by $\frac{1}{m}\sum_{j=1}^{m}e^{-l^{(t)}_{j}}-e^{-l^{(t)}_{i}}$. To smooth optimization, we introduce $\pi^{(t)}_{i}(1-\pi^{(t)}_{i})$ to control the learning rate.
Finally, we skip the $k$ layers with the lowest action preference $\mathbf{s}$. The search process can find out the redundancy in the original model for the downstream task By skipping the most redundant $k$ layers, we can gain the accelerated model in the downstream task.
Table 1: Comparison among EAS and other PETL methods for ScienceQA on LLaVA. The corner marks of PCETL methods represent the number of skipped modules. “Inference Time” refers to the time per sample. The best and second best results are marked in bold and underline, respectively.
Method Updated Parameters FLOPs Inference Time Subject Context Modality Grade Average
NAT SOC LAN TXT IMG NO G1-6 G7-12
Fully tuned methods
LLaVA-7B [liu2023visual] 100.00% 5018.4G (+0.0G) 2.19s 90.4 96.0 88.0 89.5 88.0 90.7 90.9 90.9 90.9
Parameter-efficient methods
LoRA [conf/iclr/HuSWALWWC22] 0.62% 5018.4G (+0.0G) 2.34s 86.5 95.1 85.9 86.1 86.6 87.0 88.7 87.2 88.1
Adapter [conf/cvpr/Sung0B22] 0.54% 5041.6G (+23.2G) 2.89s 86.9 95.5 85.8 86.4 86.3 87.7 89.5 86.5 88.4
MAM [conf/iclr/HeZMBN22] 0.54% 5041.6G (+23.2G) 2.87s 87.3 95.7 86.6 86.5 86.1 88.9 89.6 87.7 88.9
Parameter and computation efficient methods
DAS4 [wu2023parameter] 0.59% 4475.0G (-543.4G) 2.53s 87.5 95.5 84.5 87.4 87.0 86.4 89.2 87.0 88.4
[…]
Table 2: Comparison among EAS and other PETL-Based MLLMs for ScienceQA. The corner marks of PCETL methods represent the number of skipped modules. “Inference Time” refers to the time per sample. The best and second best results are marked in bold and underline, respectively.
Method Updated Parameters FLOPs Inference Time Subject Context Modality Grade Average
NAT SOC LAN TXT IMG NO G1-6 G7-12
Fully tuned methods
LLaVA-7B [liu2023visual] 100.00% 5018.4G (+3303.5G) 2.19s 90.4 96.0 88.0 89.5 88.0 90.7 90.9 90.9 90.9
Parameter-efficient methods
LaVIN-7B [luo2023cheap] 0.06% 1715.4G (+0.6G) 3.70s 89.3 94.9 85.2 88.5 87.5 88.1 90.2 88.1 89.4
LLaMA-LoRA [hu2022lora] 0.08% 1714.8G (+0.0G) 2.22s 84.1 78.0 85.7 82.8 74.1 88.6 83.8 82.2 83.2
LLaMA-Adapter [zhang2023llama] 0.03% 1839.6G (+124.7G) 4.35s 84.4 88.3 84.4 83.7 80.3 86.9 85.8 84.1 85.2
Parameter and computation efficient methods
LaVIN-DAS4 [wu2023parameter] 0.07% 1501.3G (-213.6G) 3.23s 89.0 94.6 85.1 87.9 86.5 88.4 89.7 88.1 89.2
[…]
Table 3: Comparison among EAS and other PETL methods on Slake and AID. The base model used is LLaVA-7B. “ZS” and “FT” denote “zero-shot” and “fine-tuned”, receptively. The best results of efficient tuning methods are marked in bold and the second best is underline, respectively.
Methods Updated Parameter Slake AID Average
FLOPs Accuracy (Open) Accuracy (Closed) FLOPs Accuracy (20%) Accuracy (50%) FLOPs Accuracy
Fully tuned methods
LLaVA-7B-ZS [liu2023visual] 100.00% 4904.0G (+0.0G) 28.7 56.5 4869.9G (+0.0G) 25.5 25.6 4886.9G (+0.0G) 32.6
LLaVA-7B-FT [liu2023visual] 100.00% 4904.0G (+0.0G) 80.3 85.6 4869.9G (+0.0G) 97.7 97.9 4886.9G (+0.0G) 90.4
Parameter-efficient methods
LoRA [conf/iclr/HuSWALWWC22] 0.61% 4904.0G (+0.0G) 79.2 85.6 4869.9G (+0.0G) 96.9 97.3 4886.9G (+0.0G) 89.8
Adapter [conf/cvpr/Sung0B22] 0.53% 4926.7G (+22.7G) 77.5 87.3 4892.4G (+22.5G) 96.2 96.8 4909.5G (+22.6G) 89.5
MAM [conf/iclr/HeZMBN22] 0.53% 4926.7G (+22.7G) 77.4 87.7 4892.4G (+22.5G) 96.2 97.7 4909.5G (+22.6G) 89.8
Parameter and computation efficient methods
DAS4 [wu2023parameter] 0.63% 4373.0G (-531.0G) 79.2 86.5 4342.9G (-527.0G) 96.6 97.8 4358.0G (-529.0G) 90.0
[…]
Method Updated Parameters VQA NLVR2 Flickr30K Average
FLOPs test-dev FLOPs test-P FLOPs IR/TR R@1 FLOPs Accuracy
Full tuned methods
Full Tuning 100.00% 93.2G (+0.0G) 77.4 52.9G (+0.0G) 83.1 93.2G (+0.0G) 82.2/94.3 (+0.0G) 79.8G (+0.0G) 84.3
Parameter-efficient methods
Shallow Prompt [conf/acl/LiL20] 0.09% 121.9G (+28.7G) 68.5 79.7G (+26.9G) 65.7 121.9G (+28.7G) 74.2/88.6 107.9G (+28.7G) 74.3
Deep Prompt [journals/corr/abs-2203-12119] 0.57% 99.8G (+6.5G) 70.8 58.5G (+5.6G) 72.6 99.8G (+6.5G) 78.8/89.4 86.0G (+6.2G) 77.9
LoRA [conf/iclr/HuSWALWWC22] 0.09% 93.2G (+0.0G) 74.0 52.9G (+0.0G) 78.8 93.2G (+0.0G) 79.9/92.6 79.8G (+0.0G) 81.3
Adapter [conf/cvpr/Sung0B22] 1.65% 94.9G (+1.6G) 74.7 54.3G (+1.4G) 79.9 94.9G (+1.6G) 80.4/91.9 81.3G (+1.6G) 81.7
MAM [conf/iclr/HeZMBN22] 1.11% 94.3G (+1.1G) 75.1 53.5G (+0.7G) 80.4 94.3G (+1.1G) 80.4/93.2 80.7G (+1.0G) 82.3
Parameter and computation efficient methods
DAS4 [wu2023parameter] 1.65% 82.1G (-11.2G) 74.8 47.7G (-5.1G) 80.1 82.1G (-11.6G) 80.1/91.8 70.6G (-9.2G) 81.7
[…]
@@ 5.1 Datasets and Metrics
Out-domain VL Benchmarks. For visual-language (VL) tasks, we first apply EAS to LLaVA on two out-domain datasets: Slake [conf/isbi/LiuZXMYW21], which is focused on medical question answering, and AID [journals/tgrs/XiaHHSBZZL17] which involves remote sensing image classification. Slake [conf/isbi/LiuZXMYW21] is designed for medical visual question answering, collecting images of CT, MRI, and X-ray scans across ten different parts of the human body, such as the lungs, abdomen, and brain. The dataset includes ten types of questions related to organ identification, location, and abnormality. It is split into train, val, and test sets, containing $9,849$, $2,109$, and $2,070$ examples, respectively. Depending on the nature of the answer, questions can be categorized as “open” or “closed”. open questions have answers in the form of words or phrases, while closed questions have binary answers (i.e., yes or no). AID [journals/tgrs/XiaHHSBZZL17] is a dataset used for remote sensing images classification. It contains 10,000 images across 30 different scene types. These images are collected from Google Earth and cover diverse regions worldwide. The dataset can be divided into subsets with “20%” and “50%” of samples used for training.
Common VL Benchmarks. We also evaluated EAS on several widely recognized VL benchmarks, including ScienceQA [lu2022learn], VQA2.0 [conf/cvpr/GoyalKSBP17], NLVR2 [conf/acl/SuhrZZZBA19], and Flickr30K [journals/ijcv/PlummerWCCHL17]. ScienceQA [lu2022learn] is a large-scale benchmark for science question answering, encompassing three main subjects, 26 topics, 127 categories, and 379 distinct skills. The dataset includes examples with text only as well as those with both text and images. It is split into train, val, and test sets, containing 12,726, 4,241, and 4,241 examples, respectively. Different from traditional question-answering datasets, ScienceQA requires the model to give the reasoning process and have a longer output sequence. VQA2.0 [conf/cvpr/GoyalKSBP17] is a visual question answering dataset that consists of 204,721 images, with each image linked to at least three open-ended questions. NLVR2 [conf/acl/SuhrZZZBA19] is a benchmark designed to determine whether a natural language statement is true given a pair of images, effectively framing it as a binary classification task based on joint representations of the two images. Flickr30K [journals/ijcv/PlummerWCCHL17], re-split by Karpathy et al. [journals/pami/KarpathyF17], consists of 31,000 images, each paired with five descriptive sentences. This dataset is commonly used for evaluating cross-modal retrieval tasks, where the goal is to retrieve relevant text given an image or vice versa.
@@ 5.2 Implementation Details
The base models employed in our study are LLaVA [liu2023visual] and LaVIN [luo2023cheap]. When applying EAS to LLaVA, we insert PIAs before each attention module, and set the hidden dimension of PIA to 256. We also employ PIAs for parameter-efficient adaptations, similar to [luo2023cheap, zhang2023llama]. These PIAs also have a hidden dimension of 128, and only one path for adaptation is used, i.e., $f_{d1}$ in Eq.4. As for LaVIN, following its default settings, we use LLaMA-7B [touvron2023llama] as the LLM, and adopt ViT-L [conf/iclr/DosovitskiyB0WZ21] as the image encoder. Regarding PIA, we set the hidden dimension to 32 for skipping and 8 for adaptation. We insert these PIAs before each attention module. For redundancy evaluation, we first train the subnetworks with randomly skipped modules for 5 epochs. Then we introduce redundancy evaluation in the next 5 epochs, where 3 subnetworks are sampled every 10 training steps. In addition, we also validate the proposed EAS on a classical VL pre-trained model named METER [conf/cvpr/DouXGWWWZZYP0022]. Its attention modules in a co-attention layer are considered as independent potential skipped components. The PIAs for skipping attention modules are configured with a hidden dimension of 144. The remaining settings following the defaults in DAS [wu2023parameter].
@@ 5.3.1 Quantitative Analysis
Comparison with PETL and PCETL methods. We first compare our EAS with PCETL and PETL methods, such as DAS [wu2023parameter] and Adapter [conf/cvpr/Sung0B22], on ScienceQA [lu2022learn] which requires long output sequences, of which results are given in Tab.1 and Tab.2. We can first observe that PETL methods achieve great performance, while retaining a very small number of parameters to update. However, the use of adapters also obviously slows down their inference speeds. For instance, the inference of LLaMA-Adapter in Tab.2 consumes more than $4.35$ seconds for each example. Compared with these approaches, both DAS and EAS can achieve competitive performance, while retaining parameter and computation efficiency. Speciffically, in Tab. 1, DAS can skip $4$ Transformer layers of LLaVA for better efficiency compare to PETL method MAM [conf/iclr/HeZMBN22] while retraining comparable performance, i.e $0.6\%$ accuracy while $11.8\%$ efficiency. Meanwhile, EAS-7B12 can improve the inference speed by $36.6\%$ with only $1.0\%$ accuracy drop.
In Tab.2, we examine EAS in multimodal transfer learning of LLMs and compare it with an PETL-based MLLM called LaVIN [luo2023cheap]. In practice, we replace the adapter-designs in LaVIN with DAS and our EAS. Compared to LaVIN, DAS4 significantly improves the inference efficiency by $1.14$ times with only $0.2\%$ accuracy drop. Furthermore, EAS12 can achieve simialr performance while improving the inference speeds by $2.18$ times compared to LaVIN. As can be seen, also as a PECTL method, our EAS is much superior in actual speed-up of MLLMs than DAS, while its performance retains comparable. We can observe that EAS is obviously superior in actual speed-up, while retaining similar performance. When we combine the experiment results in Tab.1 and Tab.2, we can find that the difficulty of transferring is related to the base model itself. Specifically, DAS4 has an average performance of 88.1% on LLaVA for ScienceQA, while the performance in LaVIN is 89.2%, achieving a 0.4% improvement. For the proposed EAS method, it can skip 12 layers on LaVIN without performance drop, but can only skip 8 layers on LLaVA.
In Tab.3, we examine EAS on two out-domain benchmarks, i.e., Slake [conf/isbi/LiuZXMYW21] and AID [journals/tgrs/XiaHHSBZZL17], and compare it with a set of PCETL and PETL methods. From this table, we can first observe that the generalization of MLLMs like LLaVA is very poor on this two benchmarks, which are about 42.6% and 25.5%, respectively. This result confirms our motivation that MLLMs have a greater demand in task tuning. It can be seen that either using full tuning or PETL methods, the performance can be substantially improved. Among the compared methods, the PECTL methods, i.e., DAS and our EAS, are the ones that can reduce computation overhead while achieving high transfer learning performance. In particular, the merits of our EAS are also more prominent than DAS, which can skip up to 8 MHAs with almost no performance drops to full tuning, i.e., 90.2 v.s. 90.4. Another observation is that the impact of EAS is also influenced by the downstream tasks. For instance, EAS can skip 12 layers without performance drop of LLaVA on Slake dataset, while only 10 layers can be skipped for AID dataset.
Table 5: Comparison of different redundancy evaluation candidates for EAS on ScienceQA. “Num. of Skip” refers to the number of skipped modules. “Updated Param.” refers to the scale of updated parameters. “Infer. Time” refers to the inference time.
@@
Candidate Modules Num. of Skip Updated Param. Infer. Time Average Performance
MHA 4 0.08% 2.08s 89.4
8 0.10% 1.91s 89.8
12 0.11% 1.69s 90.0
FFN 4 0.08% 2.26s 88.3
8 0.11% 2.18s 85.0
12 0.13% 2.11s 82.1
MHA or FFN 4 0.08% 2.13s 88.4
8 0.10% 2.09s 89.2
12 0.12% 1.78s 88.6
Figure 4: Comparison of the inference time and performance between EAS and DAS on ScienceQA dataset. EAS can better speed up the inference while maintaining the performance.
Akin to DAS [wu2023parameter], we also generalize our EAS to a classical and smaller VL pretrained model called METER [conf/cvpr/DouXGWWWZZYP0022] on three widely-used VL benchmarks in Tab. 4. Compared with the popular PETL methods [journals/corr/abs-2203-12119, conf/iclr/HuSWALWWC22, conf/iclr/HeZMBN22], our EAS can also maintain a small scale of updated parameters, while achieving comparable performance but better computation efficiency. For instance, compared with adapter, our EAS is slightly better in performance, i.e., 81.9 v.s. 81.7, but EAS can save up to 13.0% FLOPs. In terms of DAS, the advantages of our EAS are also obvious. For instance, retaining similar performance, EAS8 and EAS12 can reduce more computation by 8.0% and 12.0%, respectively. Overall, these experimental results well validate the generalization of EAS on common VL models.
Comparisons of Actual Inference Speeds. In Fig. 4, we further compare the inference speeds of our EAS and DAS on LaVIN with different settings. The first observation from Fig. 4 is that DAS can skip certain layers of LaVIN with limited performance drops, e.g., skipping 2-4 layers. Its efficiency gains become more obvious when skipping more than 6 Transformer layers, while the performance also decreases to a certain extent. Compared with DAS, EAS has obvious merits in both performance and efficiency. For instance, with the same number of skipped modules, EAS is consistently faster than DAS by $44.4\%$ to $51.6\%$. By skipping 8 MHAs, EAS can obtain more obvious performance gains than the default LaVIN, i.e., 89.8 v.s. 89.4, while the inference speed is about $1.92$ times faster. Overall, these results well validate the motivation and designs of EAS towards PCETL of MLLMs.
@@
Table 6: Ablation of Propagate-Information Adapter (PIA) on EAS-7B12. Here, Base refers to using only $f_{d1}$ in Eq. 4. “Num. of Skip” refers to the number of skipped modules. “Infer. Time” refers to the inference time.
Setting Updated Param. Infer. Time Average Performace
Base 0.09% 2.94s 87.1
+ Info. Exc. in Eq.4 0.11% 3.23s 90.1
+ Frozen bias in Eq.11 0.11% 2.97s 90.0
+ Re-param. 0.11% 1.69s 90.0
Table 7: Ablation of the skipping number on ScienceQA. “Updated Param.” refers to the scale of updated parameters. “Infer. Time” refers to the inference time.
Methods Updated Param. Infer. Time Average Performance
LaVIN-7B 0.06% 3.70s 89.4
EAS-7B4 0.08% 2.08s 89.4
EAS-7B8 0.10% 1.92s 89.8
EAS-7B12 0.11% 1.69s 90.0
EAS-7B16 0.13% 1.56s 88.4
@@
Ablation Study. To examine the designs of EAS, we further conduct extensive ablations in Tab.5-5.3.1. In Tab.5, we first ablate the choice of evaluation candidates for EAS on ScienceQA, including “MHA”, “FFN” and “MHA or FFN”. From this table, we can first see that skipping MHA is the best choice among three candidates. Notably, EAS can skip up to 12 MHAs with even better performance, while its efficiency is also much better than the others. In terms of FFN, its removal leads to slower inference speed than that of MHA. More importantly, when dropping more than 8 FFNs, the performance declines significantly, e.g., $6.99\%$ by skipping 12 FFNs. This result well confirms our argument about the roles of MHA and FFN in MLLMs. In terms of “MHA or FFN”, this is a suboptimal candidate for EAS, which can obtain a good trade-off between performance and efficiency, but it is still worse than “MHA”. Overall, these results well validate the motivation of EAS, i.e., not all attention is needed for MLLMs.
Table 8: The impact of different hidden dimensions in PIA for skip-connections of EAS on ScienceQA. “Hidden Dim.” refers to the number of hidden dimension. “Updated Param.” refers to the scale of updated parameters. “Infer. Time” refers to the inference time.
Methods Hidden Dim. Updated Param. Infer. Time Average Performance
EAS-7B8 8 0.04% 1.89s 87.6
16 0.05% 1.91s 88.0
32 0.08% 1.92s 89.8
64 0.22% 1.94s 90.1
Figure 5: The predictions of LLaVA-7B-EAS and LLaVA-7B-FT. The suffixes denotes the training by our EAS and full tuning (FT), respectively. The accurate explanations for the answer are highlight in green, while the logically incorrect ones in red.
Tab. 5.3.1 ablates the designs of PIA, which is the key component of EAS. Here, “Base” refers to using only one path in the downsampling, i.e., only $f_{d1}$ in Eq. 4. “Info. Exc.” refers to the information exchange in PIA, i.e., Eq. 4. “Frozen bias” refers to fixing the output values of average pooling in Eq. 11 after the first forward. “Re-param” refers to applying the re-parameter in inference. We can see that without the two-path design for information exchange (Info. Exc.), its performance declines greatly, i.e., $3.30\%$. This phenomenon shows that adding global pooling-based information propagation operations in PIA can effectively alleviate the damage caused by over-simplifying the attention mechanism. In PIA, a key step for its re-parameterization is freeze the dynamic bias term in Eq.11, i.e., “Frozen bias”, so that PIA can be re-parameterized into the closest linear layer in the first forward. As shown in Tab.5.3.1, this operation has very limited impact on performance, i.e., only -0.1. Thus, the following re-parameterization (re-param) can be further executed to achieve actual speed-up with the same performance. Overall, Tab. 5.3.1 well confirms the designs of information exchange and re-parameterization in PIA.